OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN. Which of the following is the point slope equation of a line with the slope = 4/5 and a point of (-2, 1) ?

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

same thing, m=4/5 sub that into y-y1=m(x-x1)

OpenStudy (anonymous):

x1=-2 y1=1

OpenStudy (anonymous):

y-1=4/5(x-(-2) y-1=4/5(x+2)

OpenStudy (anonymous):

Answer is B

OpenStudy (anonymous):

thank you SO much! I'm so tired but i need to finish thiss..you're amazing I have another coming up! :)